### v1

# Mixed Signals: Understanding Model Disagreement in Multimodal Empathy Detection

Abstract
Multimodal models play a key role in empathy detection, but their performance can suffer when modalities provide conflicting cues. To understand these failures, we examine cases where unimodal and multimodal predictions diverge. Using fine-tuned models for text, audio, and video, along with a gated fusion model, we find that such disagreements often reflect underlying ambiguity, as evidenced by annotator uncertainty. Our analysis shows that dominant signals in one modality can mislead fusion when unsupported by others. We also observe that humans, like models, do not consistently benefit from multimodal input. These insights position disagreement as a useful diagnostic signal for identifying challenging examples and improving empathy system robustness.
Mixed Signals: Understanding Model Disagreement in
Multimodal Empathy Detection
Maya Srikanth Columbia University ms6198@columbia.edu                        Run Chen Columbia University runchen@cs.columbia.edu                        Julia Hirschberg Columbia University julia@cs.columbia.edu
1 Introduction
Empathy recognition in human communication is a nuanced and multifaceted task, and a core component of socially intelligent communication systems Fung et al. (2016). Empathy, commonly defined as the capacity to understand and share the emotional experiences of others, encompasses both cognitive perspective-taking and affective resonance Baumeister and Vohs (2007). In human interactions, language, speech, and visual cues jointly convey emotional intent Holler and Levinson (2019). For example, a seemingly neutral utterance might be perceived as warm or concerned when accompanied by a sympathetic tone or facial expression. For AI systems, effectively interpreting these multimodal signals requires not only accurate unimodal representations but also robust integration of potentially conflicting information across modalities. Despite recent advances in multimodal emotion recognition Jabeen et al. (2021), empathy recognition remains particularly complex, as empathy often arises from subtle contextual cues that may not align across modalities Hasan et al. (2023).
Figure 1: Given classifications provided by a single modality, we identify cases where integrating additional modalities leads to a different prediction. We analyze these flips to understand when and why they occur.
Our work investigates such complexity by examining when and why multimodal models misclassify empathy compared to their unimodal counterparts. We extend methods from prior work on dataset difficulty and human–model agreement Swayamdipta et al. (2020); Saha et al. (2022), to the underexplored domain of empathy modeling. Through a combination of model-driven analysis and human-subject experiments, we pinpoint instances where unimodal and multimodal model predictions diverge, indicating conflicting cross-modal cues. By linking modality disagreement to human disagreement, we offer new insight into the limitations of current empathy modeling and highlight the value of disagreement-based analysis in socially grounded language tasks.
2 Related Work
Empathy Modeling.
Early computational work on empathy focused on generating emotionally relevant textual responses Rashkin et al. (2019); Li et al. (2019), but these approaches are inherently limited by the absence of non-verbal cues critical to empathic understanding. Recent datasets such as EmpathicStories++ (Shen et al., 2024), medic (Zhu et al., 2023), emmi (Galland et al., 2024) and Chen et al. (2024) address this limitation by incorporating speech, facial expressions, and interaction context, enabling more comprehensive modeling of empathy. These resources have motivated frameworks like pegs (Zhang et al., 2024), which integrate text and visual stickers for affective generation. Despite these advances, empathy remains difficult to model due to its reliance on subtle, often conflicting signals across modalities. Prior work has largely focused on improving fusion strategies under the assumption that modalities are complementary Zadeh et al. (2017); Tsai et al. (2019), but has paid less attention to when fusion may fail or introduce noise.
Dataset Difficulty.
Complementary lines of work have investigated data difficulty and model disagreement as tools for understanding model behavior. Swayamdipta et al. (2020) propose the dataset cartography method to identify hard or ambiguous training samples; Saha et al. (2022) demonstrate that difficult instances are also harder for both humans and models to explain; Wang et al. (2023)’s Learning-From-Disagreement (LFD) framework underscores the importance of examining disagreements between models to gain deeper, actionable insights into their behaviors.
Yet, these methods remain underexplored in empathy modeling, where ambiguity is often intrinsic. Our work bridges this gap by using model–modality disagreement to identify inherently ambiguous empathy examples—cases where fusion misleads models, and where annotators also exhibit uncertainty.
3 Experiment 1: Identifying Complex Examples from Modality Disagreement
Figure 2: Comparing predictions between unimodal (text, audio, video) and multimodal models. We highlight regions where model predictions agree (blue and yellow quadrants) and disagree (red and green quadrants).
Disagreement between models trained on different modalities can reveal challenging, nuanced, or ambiguous examples. Here, we identify and analyze such cases of disagreement in empathy classification using a multimodal English empathy speech dataset collected from Youtube Chen et al. (2024) (referred to as EmpSpeech) consisting of 1,718 manually annotated English speech segments labeled as empathetic or neutral (Appendix B).
Experimental Setup.
Examples in EmpSpeech are comprised of video segments spanning three modalities: text (transcript), audio (speech), and video.
We finetune two models per modality on the train set from EmpSpeech: RoBERTa Liu et al. (2019) and DeBERTa He et al. (2021) for text, HuBERT Hsu et al. (2021) and Wav2Vec2 Baevski et al. (2020) for audio, and VideoMAE Tong et al. (2022) and TimeSFormer Bertasius et al. (2021) for video (Appendix A.1). [^1] The hidden layer dimensions of all models we consider are similar. Then, we extract embeddings from each best-performing unimodal model (RoBERTa, HuBERT, and VideoMAE, Table 1) to train a multimodal fusion model that projects all three modality embeddings into a shared latent space (Appendix A.2).
Modality Model Accuracy F1
Text RoBERTa 0.75 0.73
DeBERTa 0.69 0.68
Audio HuBERT 0.72 0.71
Wav2Vec2 0.68 0.63
Video VideoMAE 0.77 0.77
TimesFormer 0.64 0.62
Fusion (All Modalities) 0.76 0.72
Table 1: Performance of fine-tuned models across modalities on the empathy classification task.
Results.
We evaluate all models (unimodal and multimodal) on the test split of EmpSpeech to identify disagreements, or examples where two models with varying input modalities assign different labels, highlighting cases where different modalities may carry ambiguous, conflicting, or modality-specific signals.
Text disagrees most frequently with the other modalities (Table 2), suggesting that it does not frequently align with speech and visual cues. In contrast, audio and video are better aligned, likely due to a shared reliance on nonverbal expressive signaling such as prosody and facial expression. Our fusion model disagrees least with text, suggesting it relies more heavily on verbal content. This also may reveal an over-reliance on text-based signals by the original annotators.
Figure 2 visualizes disagreement regions between each unimodal model and the fusion model. We plot unimodal confidence (x-axis) against fusion confidence (y-axis) in the correct label; hence confidence greater than 0.5 resulted in a correct prediction. This yields four quadrants: green (multimodal correct, unimodal incorrect), red (multimodal incorrect, unimodal correct), blue (both correct), and yellow (both incorrect). Red and green quadrants are disagreement regions which we explore to identify complex examples.
Modality Text Audio Video
Text 0.338 0.318
Audio 0.338 0.253
Video 0.318 0.253
Full 0.214 0.383 0.331
Table 2: Pairwise disagreement rates between top unimodal models and between each unimodal model and the full model, measured as the proportion of test examples with differing predictions.
Feature Red vs. Blue Green vs. Blue
p-value Direction p-value Direction
valence 0.0047 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.5166 $\mu_{\text{green}}>\mu_{\text{blue}}$
arousal 0.0065 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.0136 $\mu_{\text{blue}}>\mu_{\text{green}}$
Mean Pitch 0.0100 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.0001 $\mu_{\text{blue}}>\mu_{\text{green}}$
dominance 0.0108 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.0667 $\mu_{\text{blue}}>\mu_{\text{green}}$
Min Pitch 0.0333 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.0001 $\mu_{\text{blue}}>\mu_{\text{green}}$
Jitter 0.0347 $\mu_{\text{red}}>\mu_{\text{blue}}$ 0.0667 $\mu_{\text{green}}>\mu_{\text{blue}}$
Max Intensity 0.1260 $\mu_{\text{red}}>\mu_{\text{blue}}$ 0.0023 $\mu_{\text{green}}>\mu_{\text{blue}}$
Table 3: Statistically significant t-test results comparing red vs. blue and green vs. blue examples for audio features. See Appendix D for full table.
3.1 Modality-Based Feature Analysis
To better understand examples in disagreement regions, we extract and analyze modality-based human interpretable features.
Audio.
We extract 12 prosodic and paralinguistic features from audio signals: 9 low-level acoustic features using Praat Boersma and Weenink (1992–2022) and Parselmouth Jadoul et al. (2018), and 3 high-level affective dimensions—valence, arousal, and dominance—using a finetuned Wav2Vec2 Wagner et al. (2023). We compare feature distributions using t-tests for examples in disagreement quadrants (red and green) compared to those in the blue quadrant, signifying a non-ambiguous, easy examples. Blue examples have several significantly elevated pitch-related values than red examples (Table 3), suggesting that stronger prosodic fluctuations are frequently corroborated by other modalities. Examples in the green quadrant show significantly higher Max Intensity than in blue, potentially reflecting the role of volume-based emphasis in aiding unimodal predictions. Furthermore, affective dimensions such as valence, arousal, and dominance are significantly lower in red examples, reinforcing the idea that red examples are not simply noisy, but structurally ambiguous: they express strong unimodal signals that are complicated when analyzed alongside other modalities.
AU p (R vs B) Dir p (G vs B) Direction
AU04 0.0106 red $>$ blue 0.3682 green $>$ blue
AU12 0.0174 blue $>$ red 0.8977 green $>$ blue
AU05 0.1837 blue $>$ red <0.0001 blue $>$ green
Table 4: Statistically significant t-test results comparing AU activation rates between red vs. blue and green vs. blue. See Appendix D for full table
Video.
We examine facial action unit (AU) activations Baltrušaitis et al. (2016) from video. AU04 (Brow Lowerer), AU12 (Lip Corner Puller), and AU05 (Upper Lid Raiser) show significant differences across example types, revealing how specific facial expressions contribute to perceptual ambiguity (Table 4). AU04 is more active in red examples than blue, indicating that despite its visually strong presence, its signal conflicts with other modalities. In contrast, AU12, associated with positive affect, and AU05, which is linked to attentiveness (Friesen and Ekman, 1978), both show greater activation in blue examples than in red and green, respectively, suggesting that these expressions may serve as clearer cues that are more consistently interpreted across modalities. Our findings indicate that fine-grained facial signals may contribute to perceptual complexity in the visual stream.
Text.
Finally, visualizing umap Sainburg et al. (2021) projections of text embeddings (Figure 3) reveals that examples in disagreement regions (red and green) tend to cluster along the boundary between consistently correct (blue) and consistently incorrect (yellow) examples. Rather than forming distinct or isolated groups, disagreement examples appear in transition zones within the embedding space—areas where semantic cues are less definitive, supporting our hypothesis that red and green examples are inherently ambiguous and difficult and illustrating that modality disagreement is a reliable signal of challenging examples in empathy detection.
Figure 3: UMAP projections of text-only embeddings for empathetic (left) and neutral (right) examples, colored by modality disagreement class. Red and green points tend to cluster near the decision boundary, indicating high ambiguity.
4 Experiment 2: Characterizing Complex Examples
We further assess whether model disagreements stem from data ambiguity by conducting a human annotation study to understand whether examples where models disagree are similarly challenging for human annotators.
Annotation Setup.
We sample 204 examples evenly split across the four quadrants of each Figure 2 modality plot. For each example, annotators provide a binary judgment (empathetic or neutral) from a unimodal signal, then a judgment from the full multimodal version (see Appendix C for instructions), allowing us to track how human predictions shift with additional modality signals and understand the cognitive burden of multimodal integration. All examples were annotated by one author and one external annotator.
Quadrant Unimodal Judgment Multimodal Judgment $\Delta$
Red 0.301 0.164 -0.137
Blue 0.379 0.646 0.267
Yellow 0.225 0.329 0.104
Green 0.482 0.218 -0.264
Table 5: Cohen’s Kappa between internal and external annotators, computed separately for each quadrant and prediction round.
Results.
Annotator disagreement, measured with Cohen’s Kappa Cohen (1960), can signal complex phenomena in examples Jiang and de Marneffe (2022); Pavlick and Kwiatkowski (2019) such as uncertainty in meaning leading to discrepancies in reasoning. In disagreement regions (red and green), we see a decrease in annotator agreement between unimodal and multimodal judgments (Table 5), indicating that humans diverge when weighing signals across modalities. In contrast, annotator agreement improves on examples where unimodal and multimodal model predictions are in agreement, supporting our these examples are relatively unambiguous and reliably interpreted once the full context is available (Table 5). Collectively, these results corroborate our hypothesis that modality disagreement can serve as a valuable signal for identifying ambiguous, challenging, or complex instances that are also difficult for human annotators.
5 Discussion and Conclusion
We have demonstrated how disagreement, both between modalities and between humans and models, can serve as a diagnostic lens to understand the complexity of multimodal empathy detection, challenging the assumption that more signal from other modalities reliably yields better performance. Our analysis reveals that disagreement between unimodal and multimodal models is often not arbitrary, but instead marks the presence of subtle, ambiguous, or context-sensitive cues that challenge fusion models and human annotators alike. Our findings emphasize the necessity of high-quality, context-sensitive annotation in socially complex tasks like empathy detection, where model errors may reflect genuine human uncertainty. This framework provides a scalable method for identifying ambiguity and enhancing model reliability, especially in recognizing complex emotional states that involve inherent disagreement and uncertainty. Our work lays the foundation for several directions of future work, such as creating adversarial test sets to evaluate empathy detection systems in realistic scenarios or the identification of challenging examples for human annotation in an active learning setup to improve model robustness.
Limitations
We acknowledge several limitations in our study. Our analyses are based on a limited dataset and a small number of human annotators. Given that empathy is inherently subjective, annotations may vary due to individual interpretations, potentially introducing biases rather than reflecting universal properties of the data. Additionally, we rely on a single dataset, and future work should investigate whether the patterns we observe hold across other datasets and domains.
Our data is also derived from U.S.-based, English-language television and interview content. As such, the generalizability of our findings to multilingual or culturally diverse settings may be limited. Future research should investigate these patterns in varied cultural and linguistic environments to better assess the broader applicability of our conclusions.
Ethics Statement
We use a publicly available dataset and strictly use open-source models for analysis.
All annotations were conducted by an author and an individual affiliated with the research team. No participants were recruited via crowdsourcing or external platforms, and no monetary compensation was provided, as the annotators were contributing in a research capacity. We provide detailed information on what we ask the annotators to annotate and how we plan to use the data. The annotators willingly agreed to participate with full knowledge of the task. No sensitive or identifying information were collected from annotators.
We note that empathy expression may vary across cultures, and our findings may not generalize to non-English or non-Western contexts. We encourage future work to explore these questions in more diverse settings.
We will release all code and experimental resources at https://github.com/mayasrikm/multimodal-empathy-disagreement to support reproducibility.
Appendix A Model Training Details
Data was split into train, test and validation sets using random sampling, with an 80-10-10 split. We run fine-tuning and inference for all open-source models on an A100 GPU in Google Colab.
A.1 Unimodal Model Training Details
Each model is trained on a binary empathy classification task using precomputed 768-dimensional embeddings. We freeze all but the final two transformer layers and train for 15 epochs with a learning rate of 5e-6 and batch size of 8.
A.2 Fusion Model Details
Each unimodal model representation is independently gated and passed through an attention mechanism that computes modality-specific weights. The weighted embeddings are aggregated and classified using a three-layer feedforward network with max pooling. The fusion model is trained for 10 epochs using a learning rate of 1e-4 and includes modality dropout during training.
Appendix B Dataset Details
We use a multimodal empathy dataset Chen et al. (2024) consisting of 346 English-language videos totaling approximately 53 hours, collected from YouTube between 2020 and 2022 using keywords like “empathy” and “empathetic training.” The dataset includes empathy training sessions, therapy roleplays, interviews, TED Talks, and TV/movie scenes, comprising both acted (62%) and spontaneous (38%) speech. Each video was labeled by at least three expert annotators as either empathetic or neutral, with final labels determined by majority vote. Metadata such as speaker gender, topic, and emotional context was manually annotated, covering themes like therapy, parenting, workplace dynamics, and social relationships. From this collection, a subset of 65 videos was transcribed, diarized, and manually re-aligned using Praat to ensure accurate speaker segmentation and time alignment. This process resulted in 1,718 annotated segments with speaker labels, timestamps, transcripts, and empathy stage annotations, enabling fine-grained analysis of empathy in naturalistic and semi-scripted settings.
Appendix C Annotation Instructions
We employed two annotators, one of the paper’s authors and an non-author, both fluent English speakers based in the United States. No additional demographic information was collected, as the annotation was conducted internally for research purposes.
Annotators were asked to provide two judgments per example, labeling each as either empathetic or neutral (Figure 4). A excerpt describing empathy (drawn from the Encyclopedia of Social Psychology, Volume 1, Baumeister and Vohs (2007)) was provided to ensure a consistent conceptual foundation for annotation:
Empathy is often defined as understanding another person’s experience by imagining oneself in that other person’s situation: One understands the other person’s experience as if it were being experienced by the self, but without the self actually experiencing it. There are three commonly studied components of emotional empathy. The first is feeling the same emotion as another person (sometimes attributed to emotional contagion, e.g., unconsciously “catching” someone else’s tears and feeling sad oneself). The second component, personal distress, refers to one’s own feelings of distress in response to perceiving another’s plight. The third emotional component, feeling compassion for another person, is the one most frequently associated with the study of empathy. Cognitive empathy refers to the extent to which we perceive or have evidence that we have successfully guessed someone else’s thoughts and feelings.
Annotators were given an annotation flag indicating which modality to use for the first pass; for instance, if the flag was text, only the transcript was to be used to make the first prediction. After submitting the first judgment, annotators were then given access to the full video, including all available audio, visual, and textual information. They were then asked to provide a second prediction.
Figure 4: Annotation interface
Appendix D Full Feature Comparisons
Tables 6, 7 and 8 provide additional results from the t-tests comparing examples across different confidence quadrants. Table 6 provides an internal comparison between the disagreement quadrants. Table 7 presents the full version of the audio feature comparisons summarized in Table 3. Table 8 expands on the facial feature comparisons shown in Table 4.
Feature t-stat p-value Mean Comparison
Mean Pitch 2.453 0.0159 $\mu_{\text{red}}>\mu_{\text{green}}$
Max Intensity -2.124 0.0366 $\mu_{\text{green}}>\mu_{\text{red}}$
Max Pitch 2.016 0.0465 $\mu_{\text{red}}>\mu_{\text{green}}$
Min Pitch 2.007 0.0475 $\mu_{\text{red}}>\mu_{\text{green}}$
valence -1.908 0.0593 $\mu_{\text{green}}>\mu_{\text{red}}$
arousal 1.827 0.0705 $\mu_{\text{red}}>\mu_{\text{green}}$
speaking_rate 1.773 0.0807 $\mu_{\text{red}}>\mu_{\text{green}}$
dominance 1.712 0.0899 $\mu_{\text{red}}>\mu_{\text{green}}$
Shimmer 0.773 0.4416 $\mu_{\text{red}}>\mu_{\text{green}}$
Jitter 0.622 0.5355 $\mu_{\text{red}}>\mu_{\text{green}}$
Mean Intensity 0.544 0.5886 $\mu_{\text{red}}>\mu_{\text{green}}$
HNR 0.508 0.6129 $\mu_{\text{red}}>\mu_{\text{green}}$
Min Intensity -0.429 0.6685 $\mu_{\text{green}}>\mu_{\text{red}}$
Table 6: T-test results comparing audio features between red and green examples. Statistically significant results are bolded.
Feature p (Red vs Blue) Direction p (Green vs Blue) Direction
valence 0.0047 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.5166 $\mu_{\text{green}}>\mu_{\text{blue}}$
arousal 0.0065 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.0136 $\mu_{\text{blue}}>\mu_{\text{green}}$
Mean Pitch 0.0100 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.0001 $\mu_{\text{blue}}>\mu_{\text{green}}$
dominance 0.0108 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.0667 $\mu_{\text{blue}}>\mu_{\text{green}}$
Min Pitch 0.0333 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.0001 $\mu_{\text{blue}}>\mu_{\text{green}}$
Jitter 0.0347 $\mu_{\text{red}}>\mu_{\text{blue}}$ 0.0667 $\mu_{\text{green}}>\mu_{\text{blue}}$
Max Intensity 0.1260 $\mu_{\text{red}}>\mu_{\text{blue}}$ 0.0023 $\mu_{\text{green}}>\mu_{\text{blue}}$
Mean Intensity 0.1599 $\mu_{\text{red}}>\mu_{\text{blue}}$ 0.5329 $\mu_{\text{blue}}>\mu_{\text{green}}$
HNR 0.2217 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.2055 $\mu_{\text{blue}}>\mu_{\text{green}}$
speaking_rate 0.2723 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.9991 $\mu_{\text{green}}>\mu_{\text{blue}}$
Shimmer 0.4122 $\mu_{\text{red}}>\mu_{\text{blue}}$ 0.1541 $\mu_{\text{blue}}>\mu_{\text{green}}$
Max Pitch 0.6845 $\mu_{\text{red}}>\mu_{\text{blue}}$ 0.2647 $\mu_{\text{blue}}>\mu_{\text{green}}$
Min Intensity 0.7999 $\mu_{\text{blue}}>\mu_{\text{red}}$ 0.1571 $\mu_{\text{blue}}>\mu_{\text{green}}$
Table 7: T-test results comparing audio features between red vs. blue and green vs. blue examples. Statistically significant p-values are bolded.
AU p (Red vs Blue) Direction p (Green vs Blue) Direction
AU04: Brow Lowerer 0.0106 red $>$ blue 0.3682 green $>$ blue
AU12: Lip Corner Puller 0.0174 blue $>$ red 0.8977 green $>$ blue
AU05: Upper Lid Raiser 0.1837 blue $>$ red <0.0001 blue $>$ green
AU17: Chin Raiser 0.2256 red $>$ blue 0.9802 blue $>$ green
AU10: Upper Lip Raiser 0.2275 blue $>$ red 0.6700 green $>$ blue
AU45: Blink 0.3200 blue $>$ red 0.7462 green $>$ blue
AU07: Lid Tightener 0.3252 blue $>$ red 0.9318 blue $>$ green
AU14: Dimpler 0.4593 red $>$ blue 0.0652 green $>$ blue
AU20: Lip Stretcher 0.5701 blue $>$ red 0.7907 blue $>$ green
AU09: Nose Wrinkler 0.6211 blue $>$ red 0.7639 green $>$ blue
AU25: Lips Part 0.6227 blue $>$ red 0.7492 blue $>$ green
AU01: Inner Brow Raiser 0.6529 blue $>$ red 0.4674 green $>$ blue
AU23: Lip Tightener 0.6630 red $>$ blue 0.3474 green $>$ blue
AU28: Lip Suck 0.6735 red $>$ blue 0.9846 green $>$ blue
AU26: Jaw Drop 0.6851 red $>$ blue 0.4596 blue $>$ green
AU06: Cheek Raiser 0.7097 blue $>$ red 0.3201 green $>$ blue
AU15: Lip Corner Depressor 0.9528 red $>$ blue 0.4834 green $>$ blue
AU02: Outer Brow Raiser 0.9647 blue $>$ red 0.6677 green $>$ blue
Table 8: T-test results comparing AU activation rates between red vs. blue and green vs. blue. Bolded p-values are statistically significant.
Appendix E Feature Distributions
Figures 5 and 6 visualize the distributions of key features across confidence quadrants. Figure 5 presents the distribution of selected audio features (e.g., pitch, intensity) for red, green, and blue examples, highlighting acoustic patterns associated with model disagreement. Figure 6 shows activation rates for facial Action Units (AUs) in red, green, and blue examples, illustrating how specific facial expressions vary across agreement conditions. These visualizations complement the statistical comparisons reported in Tables 7 and 8, providing a more interpretable view of the underlying feature dynamics.
Figure 5: Distribution of audio features for red, green and blue examples across the confidence quadrants. Red examples are those correctly classified by the unimodal audio model but misclassified by the multimodal model; green examples represent the reverse. Blue examples represent those correctly classified by both the unimodal audio model and the multimodal model. Significant differences appear in pitch and intensity-based features.
Figure 6: AU activation rates for red, green, and blue examples. Red bars indicate examples where the unimodal visual model predicted correctly but the multimodal model did not (Red: Unimodal $>$ 0.5, Multimodal $<$ 0.5). Green bars show the reverse. Blue bars indicate examples where both the unimodal and multimodal models correctly predicted the label.